OpenStudy (anonymous):

A cylinder and a rectangular prism have the same volume and same height. The base of the prism is a square with side length 9 cm. What is the approximate radius of the cylinder? A. 4.5 cm B. 5.1 cm C. 12.9 cm D. 25.8 cm

OpenStudy (anonymous):

If they have both the same volume and the same height, then they must also have the same area (base.)

OpenStudy (anonymous):

This is because the volume of a cylinder or prism is the area of the base times the height.

OpenStudy (anonymous):

The area of a square is the side length, squared.

OpenStudy (anonymous):

b

OpenStudy (anonymous):

So the area of the square base of the rectangular prism is 81 cm^2.

OpenStudy (anonymous):

The area of a circle is pi * radius^2.

OpenStudy (anonymous):

pi*radius^2=81 radius^2=81/pi radius=sqrt(81/pi).

OpenStudy (anonymous):

This is approximately 5.07770625193. So the answer, indeed, is (B).

OpenStudy (anonymous):

A cylindrical can of baked beans has a diameter of 8 cm and a height of 9 cm. What is the exact volume of the can? A. 72π cm³ B. 144π cm³ C. 648π cm³ D. 2304π cm³

OpenStudy (anonymous):

The volume of a cylinder is the area times the height. The area of a circle is the radius squared times pi.

OpenStudy (anonymous):

If the diameter is 8, then the radius is 4.

OpenStudy (anonymous):

4^2=16

OpenStudy (anonymous):

The area of the circular base is pi*16.

OpenStudy (anonymous):

The height is 9, so pi*16*9=144pi.

OpenStudy (anonymous):

The answer is (B).

OpenStudy (anonymous):

Do you need any more help?